# Jailbreak Attacks and Defenses against Multimodal Generative Models: A Survey

### v1

@@ A.1 Input-level Attack
A.2 Output-level Attack
In Fig. 4, attackers focus on querying outputs across multiple input variants, systematically analyzing the model’s responses to diverse inputs. Driven by specific adversarial goals, attackers employ estimation-based and search-based attack techniques to iteratively refine these input variants.
Estimation-based Attack. Estimation-based methods typically involve approximating gradients through zeroth-order estimation techniques, where attackers rely on black-box queries to estimate the model’s gradient by observing output variations in response to small input perturbations.
$\bullet$ Text → Image (T2I) Models. DiffZOO[75] addresses the challenges of gradient unavailability and discrete prompt optimization by employing zeroth order optimization to estimate gradients. During the optimization process, DiffZOO identifies critical tokens within prompts that significantly impact model behavior and selectively replaces them to enhance the attack’s effectiveness.
@@ A.2 Output-level Attack
$\bullet$ Text → Image (T2I) Models. SneakyPrompt [76] is an automated attack framework that leverages reinforcement learning to assist in replacing tokens within the search space. Reinforcement learning is used to guide the search process, optimizing adversarial jailbreak prompts by iteratively refining them based on reward signals. The framework’s reward model is grounded in cosine similarity between adversarial and malicious images in the CLIP embedding space, effectively steering the optimization toward generating unsafe content.
III-B White-box and Gray-box Jailbreak
In a white-box attack scenario, attackers have access to the internal architecture of the target model, including encoder and generator modules. This level of access allows attackers to leverage gradient information and the model’s intermediate representations, facilitating precise adversarial modifications. In certain scenarios, attackers possess partial knowledge of the model, such as its output distribution or structural details. This type of attack is referred to as a gray-box attack.
Figure 5: Illustration of white-box jailbreak attacks against multimodal generative models at the encoder level where attackers can directly exploit the vulnerabilities within the encoder architecture.
@@ B.1 Encoder-level Attack
For encoder-level attacks, attackers are restricted to accessing only the encoders to provoke harmful responses. In this case, attackers typically seek to maximize cosine similarity within the latent space, ensuring the adversarial input retains similar semantics to the target malicious content while still being classified as safe. This adversarial objective is formalized as:
$\underset{\mathcal{X}_{adv}}{\mathrm{arg}\max}\,\,Cos\left(E_{M}\left(\mathcal{X}_{adv}\right),E_{M}\left(\mathcal{X}_{mal}\right)\right).$ (3)
@@
$\bullet$ Image + Text → Text (IT2T) Models. Jailbreak in Pieces [18] and Redteaming Attack [77] both propose to construct adversarial images paired with textual prompts to disrupt model alignment. These adversarial images are optimized through gradient-based attacks within the latent space of the encoder, effectively mapping adversarial embeddings near regions associated with harmful triggers. Notably, Jailbreak in Pieces [18] targets the CLIP [43] as the surrogate encoder, which shares the same model architecture as the built-in encoder in MLLMs.
$\bullet$ Text → Image (T2I) Models. For T2I models, both MMA-Diffusion [19] and JPA [78] employ gradient-driven techniques to iteratively optimize adversarial prompts, aligning their embeddings closely with those of the target harmful content. Additionally, MMA-Diffusion [19] applies word regularization methods to remove any sensitive terms, while JPA [93] merges inappropriate concepts into the target content embedding to counter concept removal methods.
Search-based Attack (Gray-box). In certain scenarios, attackers may lack full white-box gradient access but retain access to model output distributions, such as encoder-generated embeddings. These embeddings can be exploited to construct adversarial objectives, with search-based methods used for optimization to avoid explicit gradient requirements.
$\bullet$ Text → Image (T2I) Models. Ring-A-Bell [79] iteratively explores the latent space, adjusting adversarial inputs to align their embeddings with problematic prompts. To achieve this, the method employs a genetic algorithm [94] to optimize input modifications, enabling a guided exploration of semantic similarity. To expand the search space, RT-Attack [80] aims to find the adversarial prompt within a vast vocabulary codebook. This method utilizes a naive random search strategy guided by a two-stage adversarial objective – first optimizing semantic similarity between the adversarial and target malicious prompts, then maximizing the similarity between images generated by two prompts.
Figure 6: Illustration of white-box/gray-box jailbreak attacks against multimodal generative models at the generator level where attackers have full access to the entire model architecture.
B.4 Generator-level Attack
@@
Gradient-based Attack. Here gradient-based attacks are driven by defined adversarial optimization objectives by using gradients derived from the full generative models to refine adversarial inputs. Based on the source of supervisory signals embedded within the optimization objective, gradient-based attacks can be categorized into three types: target response-based, auxiliary generative model-based, and auxiliary classifier-based.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Most existing methods [81, 82, 83, 84, 85, 86, 87, 88, 89] operate gradient-based attacks against MLLMs typically are driven by the target responses, by maximizing the likelihood loss to achieve targeted modifications in the output probability distribution. This is formally expressed as:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\;-\sum_{i=1}^{m}{\log\left(p_{\theta}\left(\mathcal{Y}_{i}|\mathcal{X}_{adv}\right)\right)},$ (4)
@@
where $\left\{\mathcal{Y}_{i}\right\}_{i=1}^{m}$ represents the target output sequence, $m$ is the sequence length and $p_{\theta}\left(\cdot\right)$ represents the probability of next token prediction.
@@
Specifically, Image Hijacks [81] crafts prompt-matching adversarial images to match malicious behaviors to arbitrary text prompts. VisualAdv [82] reveals that a single visual adversarial example generated using a “few-shot” toxic corpus can universally jailbreak aligned models. In addition to adversarial noise as a visual component, HADES [86] introduces two additional elements: images created from harmful text typography and malicious images generated using stable diffusion [10]. Moreover, Agent Smith [87] addresses the issue of “infectious jailbreak” within a multi-agent environment. By introducing adversarial images into the memory of randomly selected agents, this method can induce the agents to generate harmful responses. In contrast to methods that primarily target the visual modality, UMK[88] and Bi-Modal Adversarial Prompt Attack (BAP) [89] conduct simultaneous adversarial attacks across both image and text modalities.
@@
$\bullet$ Text → Image (T2I) Models. Given that victim T2I models $\mathcal{M}_{\theta_{V}}$ possess safety mechanisms that can remove the sensitive concept, existing methods [90, 91, 92] often perform gradient-based adversarial attacks aiming to minimize the noise prediction loss. For instance, Prompting4Debugging (P4D) [90] generates an image with an inappropriate concept by aligning noise predictions between victim T2I models $\mathcal{M}_{\theta_{V}}$ and unconstrained T2I models $\mathcal{M}_{\theta_{U}}$, which can be formulated as:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\,\,\left\|\epsilon_{\theta_{U}}\left(z_{t}|\mathcal{X}_{mal}\right)-\epsilon_{\theta_{V}}\left(z_{t}|\mathcal{X}_{adv}\right)\right\|_{2}^{2}.$ (5)
@@
Since two separate diffusion processes will increase the computational burden, UnlearnDiffAtk [91] prepare a target image $I_{tgt}$ containing sensitive content as optimized guidance to minimize discrepancies in noise predictions by:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\,\,\left\|\epsilon-\epsilon_{\theta_{V}}\left(I_{tgt,t}|\mathcal{X}_{adv}\right)\right\|_{2}^{2}.$ (6)
@@
Similarly, by simultaneously introducing target images and auxiliary models, VA3 [92] explores adversarial attacks in an online scenario where attackers iteratively generate prompts to increase the likelihood of copyright infringement. To enhance attack success, the proposed framework utilizes an amplification strategy and adversarial prompt optimization, leveraging a multi-armed bandit approach for prompt selection.
@@
$\bullet$ Text + Image → Image (IT2I) Models. Apart from using auxiliary generative models or target responses as guidance, MMA-Diffusion [19] employs an external image classifier to produce classification supervision. Specifically, MMA-Diffusion [19] launches adversarial attacks on image editing tasks and dynamically optimizes the gradients of loss items that exceed the victim safety checker’s $\mathcal{F}_{sc}$ thresholds $T$, the method aims to minimally alter image features while bypassing the safety mechanism. This adversarial objective is formalized as:
$\mathop{\mathrm{arg}\min}\limits_{\mathcal{X}_{adv}}\,\,1_{\left\{Cos\left(f_{adv},f_{mal}\right)>T\right\}}Cos\left(f_{adv},f_{mal}\right),$ (7)
$f_{adv}=\mathcal{F}_{sc}\left(\mathcal{Y}_{adv}\right),\,\,f_{mal}=\mathcal{F}_{sc}\left(\mathcal{Y}_{mal}\right).$ (8)
IV Jailbreak Defense
In this section, we introduce current efforts made in the jailbreak defense of multimodal generative models, which includes two lines of work: Discriminative defense and Transformative defense (as shown in TABLE IV). In a discriminative setting, the defense is constrained to classification tasks for assigning binary labels. In contrast, transformative defense extends beyond classification to influencing the model’s generative process. Defenders aim to produce appropriate and safe responses in the presence of malicious or adversarial inputs.
IV-A Discriminative Defense
@@
CIDER [97] [arXiv’24] Discriminative Encoder Embedding-based I + T → T Identify similarity differences between harmful text and denoising adversarial images.
Latent Guard [98] [ECCV’24] Discriminative Encoder Embedding-based T → I Learn a latent space for detecting harmful concepts in the input text embeddings.
GuardT2I [99] [NeurIPS’24] Discriminative Encoder Embedding-based T → I First use generative models to interpret intentions behind adversarial latent embeddings.
NEARSIDE [100] [arXiv’24] Discriminative Generator Embedding-based I + T → T Obtain attack direction based on benign and adversarial embeddings for classification.
JailGuard [101] [Arxiv’23] Discriminative Output Query-based I + T → T Mutate inputs and leverage response discrepancies to detect attacks.
Espresso [102] [arXiv’24] Discriminative Output Detector-based T → I Present a content detector based on Contrastive Language-Image Pre-training.
AdaShield [103] [Arxiv’24] Transformative Input Prefix-based I + T → T Devise general defense prompts and prepend them to model inputs.
BlueSuffix [104] [Arxiv’24] Transformative Input Prefix-based I + T → T Fine-tune a generator through reinforcement learning to generate defensive suffix.
UNIGUARD [105] [Arxiv’24] Transformative Input Prefix-based I + T → T Search for additive visual noise and textual suffix to purify adversarial inputs.
POSI [106] [arXiv’24] Transformative Input Refiner-based T → I Fine-tune a LLM as an optimizer which transforms toxic prompts into clean ones.
@@
SFD [125] [arXiv’24] Transformative Generator Tuning-based T → I Align the conditional scores of “unsafe” classes or concepts with those of “safe” ones.
@@
InferAligner [131] [arXiv’24] Transformative Generator Guidance-based I + T → T Extract safety steering vectors to modify the activations of the victim model.
@@
SLD [93] [CVPR’23] Transformative Generator Guidance-based T → I Guide generation in the opposite direction of ”unsafe” prompt via classifier-free guidance.
Geom-Erasing [133] [ECCV’24] Transformative Generator Guidance-based T → I Extracts geometric information of undesire concepts that serve as negative prompts.
@@
MLLM-Protector [137] [arXiv’24] Transformative Output Refiner-based I + T → T Use detectors and detoxifiers to convert harmful responses to benign.

### v2

@@ A.2 Output-level Attack
Estimation-based Attack. To address the challenges of gradient unavailability, estimation-based methods typically rely on iterative black-box queries to estimate the model’s $\mathcal{M}_{\theta}$ gradient. This estimation is achieved by observing the deviation of optimized goals $\mathcal{L}_{opt}$ in response to small input perturbations $\delta$, which can be formulated as:
$\nabla_{x}\mathcal{L}_{opt}=\frac{\mathcal{L}_{opt}\left(\mathcal{M}_{\theta}\left(x+\delta\right)\right)-\mathcal{L}_{opt}\left(\mathcal{M}_{\theta}\left(x-\delta\right)\right)}{2\delta}.$ (3)
$\bullet$ Image + Text → Text (IT2T) Models. Based on the above formula, Zer0-Jack [75] utilizes output logits or probabilities for zeroth-order gradient estimation, reducing memory usage and query complexity by optimizing only specific image regions.
$\bullet$ Text → Image (T2I) Models. Likewise, DiffZOO [76] employs zeroth order and discrete prompt optimization to estimate gradients. During the optimization process, DiffZOO identifies critical tokens within prompts that significantly impact model behavior and selectively replaces them to enhance the attack’s effectiveness.
@@
$\bullet$ Text → Image (T2I) Models. SneakyPrompt [77] is an automated attack framework that leverages reinforcement learning to assist in replacing tokens within the search space. The framework’s reward model is grounded in cosine similarity between adversarial and malicious embeddings, effectively steering the optimization toward generating unsafe content. To expand the search space, RT-Attack [78] aims to find the adversarial prompt within a vast vocabulary codebook. This method utilizes a naive random search strategy guided by a two-stage adversarial objective – first optimizing semantic similarity between the adversarial and target malicious prompts, then maximizing the similarity between output images generated by two prompts.
Multi-turn Dialogue. Relying on the autonomous reasoning and tool-use capabilities inherent in LLMs, multi-turn dialogue methods facilitate interactions between large models and victim models, progressively intensifying malicious intent.
$\bullet$ Image + Text → Text (IT2T) Models. Some methods [79, 80] adopt self-adversarial methods to execute the jailbreak attacks. Specifically, SASP [79] employs carefully crafted dialogues to steal confidential system prompts to enhance the attack success rate. SSA [80] identifies the Safety Snowball Effect which begins with benign inputs and iteratively prompts the model for progressively more harmful outputs through context-driven interactions. Instead of relying on self-contained LLMs, IDEATOR [81] includes the additional attack models that refine prompts based on the victim model’s prior responses and integrate tool-using capabilities for malicious image generation.
$\bullet$ Text → Image (T2I) Models. Several recent works [82, 53, 84] introduce generated images as feedback for iterative refinement. For instance, APGP [82] refines high-risk prompts concerning copyright infringement based on self-generated scores which require measuring image-image consistency. CoJ [53] decomposes the query into multiple sub-queries, and then prompts T2I models to generate and iteratively edit images based on these sub-queries. Building on a collaborative multi-agent framework, Atlas [85] employs two specialized agents: the mutation agent and the selection agent. Each agent is equipped with four modules: a MLLM/LLM brain, planning, memory, and tool usage. These two agents collaborate in an iterative process to determine jailbreak prompt-triggered responses and generate new candidate jailbreak prompts.
$\bullet$ Audio → Audio (A2A) Models. Constructing targeted scenarios can induce the model to take on a role aligned with harmful behaviors. VoiceJailbreak [21] introduces a multi-step role-play attack on GPT-4o by exploiting fictional storytelling techniques. This method constructs prompts around key elements of fictional writing: setting, character, and plot. By engaging the model as a reader of fictional narratives, the adversary rephrases forbidden questions as assertive statements within fictional contexts.
III-B Gray-box and White-box Attack
In a white-box attack scenario, attackers have access to the internal architecture of the target model, including encoder and generator modules. This level of access allows attackers to leverage gradient information and the model’s intermediate representations, facilitating precise adversarial modifications.
In certain scenarios, attackers possess partial knowledge of the model, such as the architecture, some internal parameters (e.g., pre-trained encoder), or the model’s output distribution. This type of attack is referred to as a gray-box attack.
@@ B.1 Encoder-level Attack
For encoder-level attacks, attackers are restricted to accessing only the encoders to provoke harmful responses. In this case, attackers typically seek to maximize cosine similarity within the latent space, ensuring the adversarial input retains similar semantics to the target malicious content while still being classified as safe. This adversarial objective [53, 18, 86, 19, 54, 99, 100] is formalized as:
$\underset{x_{adv}}{\mathrm{arg}\max}\,\,Cos\left(E_{M}\left(x_{adv}\right),E_{M}\left(x_{mal}\right)\right).$ (4)
@@
Figure 5: Illustration of gray-box jailbreak attacks against multimodal generative models at the encoder level where attackers can directly exploit the vulnerabilities within the encoder architecture.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Jailbreak in Pieces [18] and Redteaming Attack [86] both propose to construct adversarial images paired with textual prompts to disrupt model alignment. These adversarial images are optimized through gradient-based attacks within the latent space of the encoder, effectively mapping adversarial embeddings near regions associated with harmful triggers.
$\bullet$ Text → Image (T2I) Models. Similarly, MMA-Diffusion [19] and JPA [54] both employ gradient-driven techniques to iteratively optimize adversarial prompts, aligning their embeddings closely with those of the target harmful content. Additionally, MMA-Diffusion [19] applies word regularization methods to remove any sensitive terms, while JPA [54] merges inappropriate concepts into the target content embedding to counter concept removal methods.
Search-based Attack. In certain scenarios, attackers may lack full white-box gradient access but retain access to model output distributions, such as encoder-generated embeddings. These embeddings can be exploited to construct adversarial objectives, with search-based methods used for optimization to avoid explicit gradient requirements.
$\bullet$ Text → Image (T2I) Models. Ring-A-Bell [53] iteratively explores the latent space, adjusting adversarial inputs to align their embeddings with problematic prompts. To achieve this, the method employs a genetic algorithm [101] to optimize input modifications, enabling a guided exploration of semantic similarity.
B.2 Generator-level Attack
@@
Figure 6: Illustration of white-box jailbreak attacks against multimodal generative models at the generator level where attackers have full access to the entire model architecture.
@@
Gradient-based Attack. Driven by defined adversarial goals, using gradients derived from the full generative models can precisely optimize adversarial inputs. Based on the source of supervisory signals embedded within the optimization objective, gradient-based attacks can be categorized into three types: target response-based, auxiliary generative model-based, and auxiliary classifier-based.
@@
$\bullet$ Image + Text → Text (IT2T) Models. Most methods [87, 88, 90, 89, 91, 92, 93, 94, 95] operating gradient-based attacks are typically driven by the target responses, by maximizing the likelihood loss to achieve targeted modifications in the output probability distribution. This is formally expressed as:
$\mathop{\mathrm{arg}\min}\limits_{x_{adv}}\;-\sum_{i=1}^{m}{\log\left(p_{\theta}\left(y_{i}|x_{adv}\right)\right)},$ (5)
@@
where $\left\{y_{i}\right\}_{i=1}^{m}$ represents the target output sequence, $m$ is the sequence length and $p_{\theta}\left(\cdot\right)$ represents the probability of next token prediction.
@@
Specifically, Image Hijacks [87] crafts prompt-matching adversarial images to match malicious behaviors to arbitrary text prompts. VisualAdv [88] reveals that a single visual adversarial example generated using a “few-shot” toxic corpus can universally jailbreak aligned models. In addition to adversarial noise as a visual component, HADES [92] introduces two additional elements: images created from harmful text typography and malicious images generated using stable diffusion [10]. Moreover, Agent Smith [93] addresses the issue of “infectious jailbreak” within a multi-agent environment. By introducing adversarial images into the memory of randomly selected agents, this method can induce the agents to generate harmful responses. In contrast to methods that primarily target the visual modality, UMK [94] and BAP [95] conduct simultaneous adversarial attacks across both image and text modalities. Notably, BAP [95] utilizes chain-of-thought (CoT) reasoning to enhance textual prompts through a feedback-iteration mechanism.
@@
$\bullet$ Text → Image (T2I) Models. Given that victim T2I models $\mathcal{M}_{\theta_{V}}$ possess safety mechanisms that can remove the sensitive concept, existing methods [96, 97, 98] often perform gradient-based adversarial attacks aiming to minimize the noise prediction loss. These attack methods typically require supervised guidance provided by auxiliary diffusion models [96, 98] or target images [97, 98].
For instance, P4D [96] generates an image with an inappropriate concept by aligning noise predictions between victim T2I models $\mathcal{M}_{\theta_{V}}$ and unconstrained T2I models $\mathcal{M}_{\theta_{U}}$, which can be formulated as:
$\mathop{\mathrm{arg}\min}\limits_{x_{adv}}\,\,\left\|\epsilon_{\theta_{U}}\left(z_{t}|x_{mal}\right)-\epsilon_{\theta_{V}}\left(z_{t}|x_{adv}\right)\right\|_{2}^{2},$ (6)
@@
where $z_{t}$ is the denoised image at timestep $t$. Since two separate diffusion processes will increase the computational burden, UnlearnDiffAtk [97] prepare a target image $I_{tgt}$ containing sensitive content as optimized guidance to minimize discrepancies in noise predictions by:
$\mathop{\mathrm{arg}\min}\limits_{x_{adv}}\,\,\left\|\epsilon-\epsilon_{\theta_{V}}\left(I_{tgt,t}|x_{adv}\right)\right\|_{2}^{2}.$ (7)
@@
Furthermore, by simultaneously introducing target images and auxiliary models, VA3 [98] explores adversarial attacks in an online scenario where attackers iteratively generate prompts to increase the likelihood of copyright infringement. To enhance attack success, the proposed framework utilizes an amplification strategy and adversarial prompt optimization, leveraging a multi-armed bandit approach for prompt selection.
@@
$\bullet$ Text + Image → Image (IT2I) Models. Apart from using auxiliary generative models or target responses as guidance, MMA-Diffusion [19] employs an external image classifier to produce supervised signals. Specifically, MMA-Diffusion launches adversarial attacks on image editing tasks and dynamically optimizes the gradients of loss items that exceed the victim safety checker’s $\mathcal{F}_{sc}$ thresholds $T$, the method aims to minimally alter image features while bypassing the safety mechanism. This adversarial objective is formalized as:
$\mathop{\mathrm{arg}\min}\limits_{x_{adv}}\,\,1_{\left\{Cos\left(f_{adv},f_{mal}\right)>T\right\}}Cos\left(f_{adv},f_{mal}\right),$ (8)
$f_{adv}=\mathcal{F}_{sc}\left(y_{adv}\right),\,\,f_{mal}=\mathcal{F}_{sc}\left(y_{mal}\right).$ (9)
$\bullet$ Text + Image → Text + Image (IT2IT) Models. Recent multimodal unified models such as Chameleon [40] tokenize all input modalities using non-differentiable functions. To facilitate gradient-based attacks, TSCO [56] introduces a 2-layer fully connected network as the tokenizer shortcut, providing backward gradients that enable continuous end-to-end optimization via this shortcut.
@@
CIDER [104] [arXiv’24] Discriminative Encoder Perturbation-based I + T → T Denoise input images to identify similarity difference.
Latent Guard [105] [ECCV’24] Discriminative Encoder Proxy-based T → I Learn a latent space for detecting harmful concepts in the input text embeddings.
GuardT2I [106] [NeurIPS’24] Discriminative Encoder Proxy-based T → I Use generative models to interpret intentions behind adversarial latent embeddings.
NEARSIDE [107] [arXiv’24] Discriminative Generator Perturbation-based I + T → T Obtain attack direction based on benign and adversarial embeddings for classification.
JailGuard [108] [Arxiv’23] Discriminative Output Perturbation-based I + T → T Mutate inputs and leverage response discrepancies to detect attacks.
Espresso [109] [arXiv’24] Discriminative Output Proxy-based T → I Present a content detector based on Contrastive Language-Image Pre-training.
AdaShield [110] [Arxiv’24] Transformative Input Perturbation-based I + T → T Devise general defense prompts and prepend them to model inputs.
UNIGUARD [111] [Arxiv’24] Transformative Input Perturbation-based I + T → T Search for additive visual noise and textual suffix to purify adversarial inputs.
BlueSuffix [112] [Arxiv’24] Transformative Input Refiner-based I + T → T Fine-tune a generator through reinforcement learning to generate defensive suffix.
POSI [113] [NAACL’24] Transformative Input Refiner-based T → I Fine-tune a LLM as an optimizer which transforms toxic prompts into clean ones.
@@
ES [116] [Arxiv’24] Transformative Encoder Refiner-based T → I Design a plug-and-play module to erase unsafe concepts from prompt embedding.
@@
SFD [126] [arXiv’24] Transformative Generator Tuning-based T → I Align the conditional scores of unsafe classes or concepts with those of safe ones.
@@
InferAligner [136] [EMNLP’24] Transformative Generator Guidance-based I + T → T Extract safety steering vectors to modify the activations of the victim model.
ASTRA [137] [arXiv’24] Transformative Generator Guidance-based I + T → T Steer models away from adversarial feature directions.
@@
SLD [139] [CVPR’23] Transformative Generator Guidance-based T → I Guide generation in the opposite direction of unsafe prompt via classifier-free guidance.
@@
CoCA [143] [COLM’24] Transformative Output Decoding-based I + T → T Calibrate the logit distribution by amplifying the impact of the safety prompt.
IMMUNE [144] [arXiv’24] Transformative Output Decoding-based I + T → T Use controlled decoding through a safe reward model.
MLLM-Protector [145] [EMNLP’24] Transformative Output Refiner-based I + T → T Use detectors and detoxifiers to convert harmful responses to benign.
ECSO [146] [ECCV’24] Transformative Output Refiner-based I + T → T Transform unsafe images into texts to activate models’ intrinsic safety mechanism.
@@
IV Jailbreak Defense
In this section, we introduce current efforts made in the jailbreak defense of multimodal generative models, which includes two lines of work: Discriminative defense and Transformative defense (as shown in TABLE IV). In a discriminative setting, the defense is constrained to classification tasks for assigning binary labels. In contrast, transformative defense extends beyond classification to influencing the model’s generative process, aiming to produce safe responses in the presence of malicious or adversarial inputs.
IV-A Discriminative Defense